OpenStudy (anonymous):

The Bangor Coat Factory found that the number of coats sold, N, varies directly as their advertising budget, A, and inversely as the price of each coat, P. The Bangor Coat Factory sold 5,600 coats when $32,000 was spent on advertising and the price of a coat was set at $40. Determine the number of coats sold when the amount spent on advertising is increased to $60,000.

OpenStudy (campbell_st):

Well for direct variation you will need A=kN so \[32000 = k \times 5600\] this results in \[k = \frac{40}{7}\] so 1/7 th of the selling price is spent on advertising.... then \[600000 = \frac{40}{7} \times N\] solve for N